OpenStudy (shootingstar28):

Which answer is an equation in point-slope form for the given point and slope? Point: (7, 4); Slope: 6 y+4=6(x−7) y−4=6(x+7) y−4=6(x−7) y−7=6(x+4)

OpenStudy (will.h):

Use the form y - y1 = m(x - x1) Where m is the slope

563blackghost (563blackghost):

\(\huge{y-y_{1}=m(x-x_{1})}\) \(\Large{(x_{1},y_{1})\rightarrow (x,y)}\) Is the point given....place it in each correct variable... \(\huge{m =slope}\)

OpenStudy (will.h):

Brilliant blackghost

563blackghost (563blackghost):

thank you Will.H :)

OpenStudy (will.h):

Sure mate :)

OpenStudy (shootingstar28):

So 6(x-7)?

OpenStudy (shootingstar28):

y – 4 = 6(x – 7)?

OpenStudy (will.h):

Indeed

563blackghost (563blackghost):

Correct :)

OpenStudy (shootingstar28):

Thank you!!!!

563blackghost (563blackghost):

np :)